# Long-time asymptotics of (1,3)-sign solitary waves for the damped nonlinear Klein-Gordon equation

### v1

@@ 4.4. Distance analysis III
Let $\vec{u}$ be a (1,3)-sign solitary waves . Furthermore, we assume that there exist $1\ll T_{1}<T_{2}\leq\infty$ such that for $T_{1}\leq t<T_{2}$
@@
By symmetry, we may assume that $\tilde{D}(T_{1})=\rho_{12}(T_{1})$. In this case, by Lemma 4.7 and continuity, we have for $T_{1}\leq t\leq T_{2}$.
$\rho_{13}>\tilde{D}+D^{\frac{1}{3}},\ \rho_{23}>\tilde{D}+D^{\frac{1}{3}}.$
@@
Here, note that by definition we have $|D_{mod}-\hat{D}|\leq c_{\heartsuit}$. We first consider the case $D_{mod}=\rho_{1}$. Since $\rho_{1}\leq\rho_{3}+c_{\heartsuit}$, by (2.11) we have
@@
Since we have
$|\rho_{1}-D|+|\rho_{2}-D|\lesssim D^{\frac{99}{100}},$
@@
If we assume $\rho_{2}>D+D^{\frac{99}{100}}$, then we have
@@
On the other hand, if $\rho_{2}<D+D^{\frac{99}{100}}$, then by Lemma 4.3 we have
$\left|c_{12}-\frac{1}{2}\right|+\left|u_{12}\cdot u_{2}-\frac{1}{2}\right|+\left|u_{1}\cdot u_{21}-\frac{1}{2}\right|\lesssim D^{-\frac{1}{100}}$
@@
The case $D_{mod}=\rho_{2}$ can be treated by the same argument as above. Next, we consider the case $D_{mod}=\rho_{3}+c_{\heartsuit}$. When $\rho_{1}$ and $\rho_{2}$ satisfy
@@
We now combine the discussion so far. First, if $c_{13}\geq 0$ or $c_{23}\geq 0$, then we have
@@
From the above, it follows that for $T_{1}\leq t\leq T_{2}$,
@@
$\tilde{D}(t^{\prime})-\hat{D}^{\prime}(t^{\prime})\leq D^{\frac{1}{5}}(t^{\prime}).$ (4.60)
@@
$|D(t)-D(t^{\prime})|$ $\lesssim D^{\frac{2}{9}}(t^{\prime}),$
$\int_{t^{\prime}}^{T_{1}}\mathcal{F}(D(s))ds$ $\sim D^{\frac{2}{9}}(t^{\prime}).$
Then, we have for $t^{\prime}\leq t\leq T_{1}$
@@
Therefore, by Lemma 4.8, we have
$\left(D_{mod}(T_{1})-\tilde{D}(T_{1})\right)-\left(D_{mod}(t^{\prime})-\tilde{D}(t^{\prime})\right)$ $\gtrsim\int_{t^{\prime}}^{T_{1}}\mathcal{F}(D(s))ds$
$\gtrsim D^{\frac{2}{9}}(t^{\prime}).$
Thus, we have
$\tilde{D}(T_{1})-\hat{D}(T_{1})\lesssim-D^{\frac{2}{9}}(t^{\prime}),$
@@ 5. A rough equilateral-triangle decomposition
By Proposition 4.1, the contributions coming from $\rho_{12},\ \rho_{23},\ \rho_{13}$ are negligible compared with $\mathcal{F}(D)$. In particular, this estimate implies that (2.19) also holds, and hence the dynamics of $z_{0},z_{1},z_{2},z_{3}$ can be described solely in terms of the equations involving $\rho_{1},\rho_{2},\rho_{3}$.
Lemma 5.1.
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then $z_{0},z_{1},z_{2},z_{3}$ satisfy for $t\gg 1$
$\begin{aligned} \dot{z}_{0}&=-\mathcal{F}(\rho_{1})u_{1}-\mathcal{F}(\rho_{2})u_{2}-\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{z}_{1}&=\mathcal{F}(\rho_{1})u_{1}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{z}_{2}&=\mathcal{F}(\rho_{2})u_{2}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{z}_{3}&=\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right).\end{aligned}$ (5.1)
Proof.
By Proposition 4.1, we have
$\mathcal{F}(\tilde{D})\lesssim e^{-D^{\frac{1}{5}}}\mathcal{F}(D)\lesssim\frac{\mathcal{F}(D)}{D^{8}}.$
In particular, (2.19) holds true. Thus, by Lemma 2.4, we obtain (5.1). ∎
In particular, using the above estimate and repeating the same computation as in Lemma 3.3, we also obtain the following lemma.
@@
    $Z_{1},Z_{2},Z_{3}$ satisfy for $t>0$,
    $\begin{aligned} \dot{Z}_{1}&=2\mathcal{F}(\rho_{1})u_{1}+\mathcal{F}(\rho_{2})u_{2}+\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{Z}_{2}&=\mathcal{F}(\rho_{1})u_{1}+2\mathcal{F}(\rho_{2})u_{2}+\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{Z}_{3}&=\mathcal{F}(\rho_{1})u_{1}+\mathcal{F}(\rho_{2})u_{2}+2\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right).\end{aligned}$ (5.2)
@@
    $\rho_{1},\ \rho_{2},\ \rho_{3}$ satisfy
    $\begin{aligned} \dot{\rho_{1}}&=2\mathcal{F}(\rho_{1})+\mathcal{F}(\rho_{2})c_{12}+\mathcal{F}(\rho_{3})c_{13}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{\rho}_{2}&=\mathcal{F}(\rho_{1})c_{12}+2\mathcal{F}(\rho_{2})+\mathcal{F}(\rho_{3})c_{23}+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right),\\ \dot{\rho}_{3}&=\mathcal{F}(\rho_{1})c_{13}+\mathcal{F}(\rho_{2})c_{23}+2\mathcal{F}(\rho_{3})+O\left(\frac{\mathcal{F}(D)}{D^{8}}\right).\end{aligned}$ (5.3)
@@
    $c_{12},c_{13},c_{23}$ satisfy
    $\begin{aligned} \dot{c}_{12}&=(1-c_{12}^{2})\left(\frac{\mathcal{F}(\rho_{1})}{\rho_{2}}+\frac{\mathcal{F}(\rho_{2})}{\rho_{1}}\right)+\mathcal{F}(\rho_{3})\left(\frac{c_{23}-c_{13}c_{12}}{\rho_{1}}+\frac{c_{13}-c_{23}c_{12}}{\rho_{2}}\right)\\ &\quad+O\left(\frac{\mathcal{F}(D)}{D^{9}}\right),\\ \dot{c}_{13}&=(1-c_{13}^{2})\left(\frac{\mathcal{F}(\rho_{1})}{\rho_{3}}+\frac{\mathcal{F}(\rho_{3})}{\rho_{1}}\right)+\mathcal{F}(\rho_{2})\left(\frac{c_{23}-c_{12}c_{13}}{\rho_{1}}+\frac{c_{12}-c_{13}c_{23}}{\rho_{3}}\right)\\ &\quad+O\left(\frac{\mathcal{F}(D)}{D^{9}}\right),\\ \dot{c}_{23}&=(1-c_{23}^{2})\left(\frac{\mathcal{F}(\rho_{2})}{\rho_{3}}+\frac{\mathcal{F}(\rho_{3})}{\rho_{2}}\right)+\mathcal{F}(\rho_{1})\left(\frac{c_{13}-c_{12}c_{23}}{\rho_{2}}+\frac{c_{12}-c_{13}c_{23}}{\rho_{3}}\right)\\ &\quad+O\left(\frac{\mathcal{F}(D)}{D^{9}}\right).\end{aligned}$ (5.4)
Using these equations, we analyze the long-time behavior of $(1,3)$-sign solitary waves.
@@ 5.1. Estimates for $D$
We first derive estimates for $D$.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have
@@
$\dot{D}(t^{\prime})\leq\frac{1}{1000}\mathcal{F}(D(t^{\prime})).$ (5.10)
By symmetry, we may assume $\rho_{1}=D$. If $\rho_{2}>D+D^{\frac{1}{2}}$ or $\rho_{3}>D+D^{\frac{1}{2}}$ holds, then
@@
which contradicts Proposition 4.1. Therefore we obtain (5.9). Then, by (5.9), we obtain (5.5) and (5.6), and we complete the proof. ∎
@@ 5.2. A Lyapunov-function argument
In this subsection, we use a Lyapunov function to show that, as $t\to\infty$, the quantities $\rho_{1},\rho_{2},\rho_{3}$ converge to specific sizes that are coupled to $c_{12},\ c_{23},\ c_{13}$. As a preparation, we define $b_{1},b_{2},b_{3}$ as
@@
We define $x=(x_{1},x_{2},x_{3})$. Then we have
@@
Therefore we obtain (5.19). Next, we prove (5.20). Since we have
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then we have for $k=1,2,3$
@@
Note that $\zeta_{k}$ are bounded, and moreover
@@
Using this, for $k=1,2,3$ we have
@@
By (5.30), $\zeta_{1}$ satisfies
@@ 5.3. Long-time behavior of the inner products
By Lemma 5.6, we can estimate $\mathcal{F}(\rho_{1}),\mathcal{F}(\rho_{2}),\mathcal{F}(\rho_{3})$ in terms of formulas involving inner products. Using these estimates, we refine the relations for $c_{12},\ c_{13},\ c_{23}$. Before stating the lemma, we define $\mathcal{N}$ as
@@
Let $\vec{u}$ be a $(1,3)$-sign solitary waves. Then, we have
@@
First, note that for $k=1,2,3$,
@@
By (5.45) and (5.46), we have
$\begin{aligned} &\quad(1+c_{12})\left((2-c_{23})(2+c_{23}-c_{12}-c_{13})+(2-c_{13})(2+c_{13}-c_{12}-c_{23})\right)\\ &\quad+(2-c_{12})(2+c_{12}-c_{23}-c_{13})(c_{13}+c_{23})\\ &=8+4c_{12}+2c_{13}+2c_{23}-4c_{12}^{2}-3c_{13}^{2}-3c_{23}^{2}+4c_{12}c_{13}c_{23}\\ &\quad-2c_{13}c_{23}-c_{12}c_{13}-c_{12}c_{23}.\\ &=\mathcal{N}-c_{12}^{2}-c_{13}c_{23}+2c_{12}.\end{aligned}$ (5.47)
By (5.43), (5.44), and (5.47), the expression for $\dot{c}_{12}$ indeed satisfies (5.39). The same computation for $\dot{c}_{13}$ and $\dot{c}_{23}$ yields the corresponding identities in (5.39), and the lemma follows. ∎

### v2

@@ 4.4. Distance analysis III
Let $\vec{u}$ be a (1,3)-sign solitary waves . Furthermore, we assume that there exist $1\ll T_{1}<T_{2}\leq\infty$ such that we have for $T_{1}\leq t<T_{2}$
@@
By symmetry, we may assume that $\tilde{D}(T_{1})=\rho_{12}(T_{1})$. In this case, by Lemma 4.7 and continuity, we have for $T_{1}\leq t<T_{2}$.
$\rho_{12}(t)=\tilde{D}(t),\ \rho_{13}(t)>\tilde{D}(t)+D^{\frac{1}{3}}(t),\ \rho_{23}(t)>\tilde{D}(t)+D^{\frac{1}{3}}(t).$
@@
We note that by definition we have $|D_{mod}-\hat{D}|\leq c_{\heartsuit}$. First, we assume that $D_{mod}=\rho_{1}$. Since we have $\rho_{1}\leq\rho_{3}+c_{\heartsuit}$, by (2.11) we obtain
@@
Since we have for $T_{1}\leq t<T_{2}$
$|\rho_{1}-D|+|\rho_{12}-D|\leq D^{\frac{99}{100}},$
$\rho_{2}-D\geq-D^{\frac{99}{100}},$
@@
Furthermore, if we have $\rho_{2}>D+D^{\frac{99}{100}}$, then we have
@@
On the other hand, if we have $\rho_{2}<D+D^{\frac{99}{100}}$, then by Lemma 4.3 we have
$\left|c_{12}-\frac{1}{2}\right|+\left|u_{12}\cdot u_{2}-\frac{1}{2}\right|+\left|u_{1}\cdot u_{21}-\frac{1}{2}\right|\lesssim D^{-\frac{1}{100}},$
@@
The case $D_{mod}=\rho_{2}$ can be treated by the same argument as above. Next, we assume that $D_{mod}=\rho_{3}+c_{\heartsuit}$. When $\rho_{1}$ and $\rho_{2}$ satisfy
@@
We now combine the arguments above. First, if $c_{13},c_{23}$ satisfy $c_{13}\geq 0$ or $c_{23}\geq 0$, then we have
@@
From the above, we have for $T_{1}\leq t<T_{2}$,
@@
$\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})\leq D^{\frac{1}{5}}(t^{\prime}).$ (4.63)
@@
$\left|D(t)-D(t^{\prime})\right|\leq 2D^{\frac{2}{9}}(t^{\prime}),$
$c_{\bullet}D^{\frac{2}{9}}(t^{\prime})\leq$ $\int_{t^{\prime}}^{T_{1}}\mathcal{F}(D(s))ds\leq\frac{1}{c_{\bullet}}D^{\frac{2}{9}}(t^{\prime}).$
Then, by Lemma 3.3, we have for $t^{\prime}\leq t\leq T_{1}$
$\tilde{D}(t)-\hat{D}(t)$ $\lesssim\tilde{D}(t^{\prime})-\hat{D}(t^{\prime})+D^{\frac{2}{9}}(t)$
$\lesssim D^{\frac{2}{9}}(t^{\prime}),$
and therefore we have for $t^{\prime}\leq t\leq T_{1}$
@@
Thus, by Lemma 4.8 and (4.51) and (4.63), there exists $c>0$ such that
$\quad\left(\hat{D}(T_{1})-\tilde{D}(T_{1})\right)+D^{\frac{1}{5}}(t^{\prime})+2c_{\heartsuit}$
$\geq\left(D_{mod}(T_{1})-\tilde{D}(T_{1})\right)-\left(D_{mod}(t^{\prime})-\tilde{D}(t^{\prime})\right)$
$\geq c\int_{t^{\prime}}^{T_{1}}\mathcal{F}(D(s))ds$
$\geq cc_{\bullet}D^{\frac{2}{9}}(t^{\prime}).$
By the above estimate, we obtain
$\tilde{D}(T_{1})-\hat{D}(T_{1})\leq-cc_{\bullet}D^{\frac{2}{9}}(t^{\prime})+D^{\frac{1}{5}}(t^{\prime})+2c_{\heartsuit},$
@@ 5. A rough equilateral-triangle decomposition
By Proposition 4.1, the contributions coming from $\rho_{12},\ \rho_{23},\ \rho_{13}$ are negligible compared with $\mathcal{F}(D)$. In particular, this estimate implies that (2.19) holds, and hence we can simplify the leading term in Corollary 3.4.
@@
    $z_{0},z_{1},z_{2},z_{3}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{z}_{0}&=-\mathcal{F}(\rho_{1})u_{1}-\mathcal{F}(\rho_{2})u_{2}-\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{z}_{1}&=\mathcal{F}(\rho_{1})u_{1}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{z}_{2}&=\mathcal{F}(\rho_{2})u_{2}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{z}_{3}&=\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right).\end{aligned}$ (5.1)
@@
    $Z_{1},Z_{2},Z_{3}$ satisfy for $t\gg 1$,
    $\begin{aligned} \dot{Z}_{1}&=2\mathcal{F}(\rho_{1})u_{1}+\mathcal{F}(\rho_{2})u_{2}+\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{Z}_{2}&=\mathcal{F}(\rho_{1})u_{1}+2\mathcal{F}(\rho_{2})u_{2}+\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{Z}_{3}&=\mathcal{F}(\rho_{1})u_{1}+\mathcal{F}(\rho_{2})u_{2}+2\mathcal{F}(\rho_{3})u_{3}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right).\end{aligned}$ (5.2)
@@
    $\rho_{1},\ \rho_{2},\ \rho_{3}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{\rho_{1}}&=2\mathcal{F}(\rho_{1})+\mathcal{F}(\rho_{2})c_{12}+\mathcal{F}(\rho_{3})c_{13}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{\rho}_{2}&=\mathcal{F}(\rho_{1})c_{12}+2\mathcal{F}(\rho_{2})+\mathcal{F}(\rho_{3})c_{23}+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right),\\ \dot{\rho}_{3}&=\mathcal{F}(\rho_{1})c_{13}+\mathcal{F}(\rho_{2})c_{23}+2\mathcal{F}(\rho_{3})+O\left(\frac{\mathcal{F}(D)}{D^{6}}\right).\end{aligned}$ (5.3)
  4. (4)
    $c_{12},c_{13},c_{23}$ satisfy for $t\gg 1$
    $\begin{aligned} \dot{c}_{12}&=(1-c_{12}^{2})\left(\frac{\mathcal{F}(\rho_{1})}{\rho_{2}}+\frac{\mathcal{F}(\rho_{2})}{\rho_{1}}\right)+\mathcal{F}(\rho_{3})\left(\frac{c_{23}-c_{13}c_{12}}{\rho_{1}}+\frac{c_{13}-c_{23}c_{12}}{\rho_{2}}\right)\\ &\quad+O\left(\frac{\mathcal{F}(D)}{D^{7}}\right),\\ \dot{c}_{13}&=(1-c_{13}^{2})\left(\frac{\mathcal{F}(\rho_{1})}{\rho_{3}}+\frac{\mathcal{F}(\rho_{3})}{\rho_{1}}\right)+\mathcal{F}(\rho_{2})\left(\frac{c_{23}-c_{12}c_{13}}{\rho_{1}}+\frac{c_{12}-c_{13}c_{23}}{\rho_{3}}\right)\\ &\quad+O\left(\frac{\mathcal{F}(D)}{D^{7}}\right),\\ \dot{c}_{23}&=(1-c_{23}^{2})\left(\frac{\mathcal{F}(\rho_{2})}{\rho_{3}}+\frac{\mathcal{F}(\rho_{3})}{\rho_{2}}\right)+\mathcal{F}(\rho_{1})\left(\frac{c_{13}-c_{12}c_{23}}{\rho_{2}}+\frac{c_{12}-c_{13}c_{23}}{\rho_{3}}\right)\\ &\quad+O\left(\frac{\mathcal{F}(D)}{D^{7}}\right).\end{aligned}$ (5.4)
Proof.
By Proposition 4.1, we have
$\mathcal{F}(\tilde{D})\lesssim e^{-D^{\frac{1}{5}}}\mathcal{F}(D)\lesssim\frac{\mathcal{F}(D)}{D^{6}}.$
In particular, (2.19) holds true. Thus, by Lemma 2.4 and Corollary 3.4, we complete the proof. ∎
Using these equations, we analyze the long-time behavior of $(1,3)$-sign solitary waves solutions.
@@ 5.1. Estimates for $D$
First, we give bounds on $D$.
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then, we have for $t\gg 1$
@@
$\dot{D}(t^{\prime})<\frac{1}{1000}\mathcal{F}(D(t^{\prime})).$ (5.10)
By symmetry, we may assume that $\rho_{1}=D$. If $\rho_{2}>D+D^{\frac{1}{2}}$ or $\rho_{3}>D+D^{\frac{1}{2}}$ holds, then
@@
which contradicts Proposition 4.1. Therefore, we obtain (5.9). Then, by (5.9), we obtain (5.5) and (5.6), and we complete the proof. ∎
@@ 5.2. A Lyapunov-function argument
In this subsection, we use a Lyapunov function to show that, as $t\to\infty$, the quantities $\rho_{1},\rho_{2},\rho_{3}$ converge to asymptotic values determined by $c_{12},\ c_{23},\ c_{13}$. As a preliminary step, we define $b_{1},b_{2},b_{3}$ as
@@
We define $x=(x_{1},x_{2},x_{3})$. Then, we have
@@
Therefore, we obtain (5.19). Next, we prove (5.20). Since we have
@@
Let $\vec{u}$ be a (1,3)-sign solitary waves. Then we have for $k=1,2,3$ and $t\gg 1$
@@
Note that $\zeta_{k}$ are bounded, and moreover we have
@@
By the above estimate and (2.11), we have for $k=1,2,3$ and $t\gg 1$
@@
By (5.26), (5.30), and
$\begin{pmatrix}2&c_{12}&c_{13}\\ c_{12}&2&c_{23}\\ c_{13}&c_{23}&2\end{pmatrix}\begin{pmatrix}b_{1}\\ b_{2}\\ b_{3}\end{pmatrix}=\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$,
$\zeta_{1}$ satisfies
@@ 5.3. Long-time behavior of the inner products
By Lemma 5.5, we may estimate $\mathcal{F}(\rho_{1}),\mathcal{F}(\rho_{2}),\mathcal{F}(\rho_{3})$ in terms of formulas involving inner products. Using these estimates, we refine the relations among $c_{12},\ c_{13},\ c_{23}$. Before stating the lemma, we define $\mathcal{N}$ as
@@
Let $\vec{u}$ be a $(1,3)$-sign solitary waves. Then, we have for $t\gg 1$
@@
We recall the definition (5.29). Then, by (5.27), we have for $k=1,2,3$
@@
By (5.45) and (5.46), we obtain
$\begin{aligned} &\quad(1+c_{12})\left((2-c_{23})(2+c_{23}-c_{12}-c_{13})+(2-c_{13})(2+c_{13}-c_{12}-c_{23})\right)\\ &\quad+(2-c_{12})(2+c_{12}-c_{23}-c_{13})(c_{13}+c_{23})\\ &=8+4c_{12}+2c_{13}+2c_{23}-4c_{12}^{2}-3c_{13}^{2}-3c_{23}^{2}+4c_{12}c_{13}c_{23}\\ &\quad-2c_{13}c_{23}-c_{12}c_{13}-c_{12}c_{23}\\ &=\mathcal{N}-c_{12}^{2}-c_{13}c_{23}+2c_{12}.\end{aligned}$ (5.47)
By (5.43), (5.44), and (5.47), $\dot{c}_{12}$ satisfies (5.39). The same computation for $\dot{c}_{13}$ and $\dot{c}_{23}$ yields the corresponding identities in (5.39), and the lemma follows. ∎